Gerhan Company's flexible budget for the units manufactured in May shows $15,600 of total factory overhead; this output level represents 70% of available capacity. During May, the company applied overhead to production at the rate of $3.00 per direct labor hour (DLH), based on a denominator volume level of 6,120 DLHs, which represents 90% of available capacity. The company worked 6,000 DLHs and incurred $18,500 of total factory overhead cost during May, including $6,800 for fixed factory overhead. Under a three-variance breakdown (decomposition) of the total overhead variance, what is the total factory overhead spending variance (to the nearest whole dollar) for May? (Round your intermediate calculation to 2 decimal places.)

Multiple Choice

  • N/A—this variance does not exist in a three-variance analysis of the total overhead variance.

  • $304 favorable.

  • $384 unfavorable.

  • $484 unfavorable.

  • $1,164 unfavorable.
Answer #1

Factory overhead variance means the variance of factory overhead and it is shown the deference in the actual and budget factory overhead during the year.

Total factory overhead spending variance = Actual factory overhead - Budgeted allowance based on Actual hours

Budgeted allowance based on Actual hours = Actual direct labor hours * Predetermined overhead rate

= 6,000 * 3 = 18,000

Actual factory overhead = 18,500

Total factory overhead spending variance = 18,500 - 18,000 = 500 (Unfavourable)